### v1

# Random Time-hopping Secure Ranging Strategy Against Distance-Reduction Attacks in UWB

Abstract
In order to mitigate the distance reduction attack in Ultra-Wide Band (UWB) ranging, this paper proposes a secure ranging scheme based on a random time-hopping mechanism without redundant signaling overhead. Additionally, a secure ranging strategy is designed for backward compatibility with existing standards such as IEEE 802.15.4a/z, combined with an attack detection scheme. The effectiveness and feasibility of the proposed strategy are demonstrated through both simulation and experimental results in the case of the Ghost Peak attack, as demonstrated by Patrick Leu et al . The random time-hopping mechanism is verified to be capable of reducing the success rate of distance reduction attacks to less than 0.01%, thereby significantly enhancing the security of UWB ranging.
Index Terms:
Ultra-Wideband, secure ranging, distance reduction attack, random time-hopping mechanism
I Introduction
With non-sine-wave narrow pulses for ranging, positioning and data transmission, Ultra-Wideband (UWB) has the characteristics of strong anti-multipath fading ability and low power consumption, and can provide centimeter-level ranging precision, which is widely used in ranging and short-range communication scenarios[1][2].
However, there always exist numerous security threats in UWB ranging. For the IEEE 802.15.4a protocol, Poturalski et al.[3] proposed the Cicada attack which continuously injects UWB pulses into the receiver during the legitimate transmission of the preamble, in addition to the Early Detection/Late Commitment (ED/LC) attack which leverages the predictability of the signal structure within the preamble[4]. To overcome known range-reduction attacks, IEEE has released a new version standard 802.15.4z[5], which enhances the precision and security of ranging by introducing Scrambled Timestamp Sequences (STS) encrypted by the Advanced Encryption Standard (AES). But in [6], Miridula Singh proposed the Cicada++ attack which executes the distance attack by transmitting pseudo-random STS signals to alter the timestamps of received signals, and further proposed the Adaptive Injection Attack (AIA) which can further refine the attack precision by controlling the placement of injected attack signals. Patrick Leu et al.[7] demonstrated the Ghost Peak attack achieving a success rate of up to 4% on commercially available Apple U1 and Qorvo UWB chips. In [8], Claudio Anliker et al. also proposed and demonstrated the Mix-Down attack, which exploits the clock drift of transceivers. These studies confirm that there are still security gaps under the new standard.
To defend against these distance attacks, there are some methods that have been proposed, e.g., a ranging scheme combining Time of Flight (TOF) and Received Signal Strength (RSS) is proposed in [9], aiming to effectively mitigate distance fraud. Additionally, Chen H et al.[10] designed the UnSpoof UWB localization system which can pinpoint the position of both the attacker and the legitimate device. And Kiseok Kim et al. [11] proposed a UWB localization system for vehicles based on Directed-Acyclic Graph (DAG) structure to enhance security. Meanwhile, effectively detecting attacks in the process of UWB security ranging is also a problem to be considered. In [12], Mridula Singh presented a novel modulation technique to detect distance enlargement attacks relying on the interleaving of pulses of different phases. Kyungho Joo et al. [13] achieved an attack detection success rate of 96.24% by leveraging the consistency of cross-correlation results between the sub-fields of STS and local templates.
Nevertheless, most of the defense schemes against distance attacks modify the established UWB physical layer standards to a large extent, increasing the costs of practical deployments. Thus based on our previous work that proposed an attack detection scheme using channel reciprocity and autoencoder [14], we further proposed a security strategy adopting a random time-hopping mechanism. By changing the synchronization time between the attack signal and the legitimate signal, distance reduction attacks can be avoided without altering the established standards. The strategy is effective against attacks that require prior knowledge by sniffing the transmitting and receiving of legitimate messages, such as Cicada++, Ghost Peak, ED/LC, Adaptive Injection, etc. And the superior performance of this strategy under the Ghost Peak attack is verified in this paper. Our main contributions can be summarized as follows:
Figure 1: Classic UWB ranging model and random time-hopping ranging model
  • A security strategy against distance attacks based on random time-hopping is proposed. The theoretical derivation of the strategy is given, and the feasibility of the scheme is verified through simulation and experiment.
  • A scheme without consuming redundant signaling to ensure that both the transmitter and the receiver are synchronized regarding the time-hopping value is proposed. By pre-storing the optional time-hopping value as a hash table and then randomly selecting it, the difficulty of subsequent attacks is increased dramatically.
II System Model and problem Statement
II-A Classic UWB Ranging Model and Attack
To mitigate the effects of factors such as clock drift, Double Side-Two Way Ranging (DS-TWR) is often used in UWB systems, which requires three ranging messages to be exchanged between both ranging sides. The distance is estimated as[5]:
$d=c\cdot T_{\text{prop}}=\frac{T_{\text{{round}}_{\text{1}}}\times T_{\text{{round}}_{\text{2}}}-T_{\text{{ reply}}_{\text{1}}}\times T_{\text{{ reply}}_{\text{2}}}}{T_{\text{{ round}}_{\text{1}}}+T_{\text{{ round}}_{\text{2}}}+T_{\text{{reply}}_{\text{1}}}+T_{\text{{ reply}}_{\text{2}}}}c,$ (1)
where the time intervals $T_{\text{{round}}_{\text{1}}}$, $T_{\text{{round}}_{\text{2}}}$, $T_{\text{{ reply}}_{\text{1}}}$, $T_{\text{{ reply}}_{\text{2}}}$ are shown in Fig. 1, $c$ refers to the speed of light.
The timestamps in the DS-TWR process in High Rate Pulse (HRP) mode[5] are obtained using a leading edge detection algorithm, which acts on the cross-correlation spectrum of the STS segments to determine the earliest arrival time of the received signal. As shown in Fig. 2, the Maximum Peak to Early Peak Ratio (MPEP), which represents the ratio between the main path and the first path power, and the Peak to Average Power Ratio (PAPR), which represents the ratio between the peak and the average power, are two decisive thresholds of the leading edge detection algorithm based on the Back-Search Time Window (BTW). The further accurate estimation of timestamps by the leading edge detection algorithm also provides attackers with more opportunities to improve the attack success probability, which is mainly exploited by many distance attacks represented by the Ghost Peak attack.
Figure 2: Cross-correlation spectrum between received STS and loacl STS
STS encrypted by AES are often considered unpredictable, which is the premise of various distance attacks. In the Ghost Peak attack as shown in Fig. 1, after the attacker sniffs the transmit time of the ranging message, it would send the attack signal during the reception of the Response message or Final message of the legitimate device. The STS segment of the attack signal is forged and the power of the segment is much higher than that of the legitimate signal. As shown in Fig. 2, the Ghost Peak attack results in a fake path in the cross-correlation spectrum. Additionally, the attack synchronization time, i.e., the time difference between the transmit time of the attack message and the corresponding legitimate message, significantly impacts the attack success rate in the Ghost Peak attack, and this will be elaborated in detail in the simulation results in Sec. IV.
II-B Random Time-hopping Ranging Model Against Attacks
For distance attacks which require the transmit time of the attack signal to be highly aligned with that of the legitimate signal, we propose a secure ranging scheme based on the random time-hopping mechanism. Similar to the frequency-hopping technique applied in spread spectrum communication systems with robust anti-interference capability, the random time-hopping mechanism ensures the security of UWB ranging by randomly transmitting messages at different time instants to increase the attack synchronization time, i.e., reducing the degree of alignment between the attack signal and the corresponding legitimate signal transmit time.
Taking the Ghost Peak attack on the Response message as an example, Fig. 1 demonstrates the anti-attack ranging model based on random time-hopping. The Responder selects a delay $\Delta t$ that can resist the attack and does not affect the whole ranging process, so that the Response message is delayed with $\Delta t$ to be transmitted. The modified $T_{\text{{round}}_{\text{1}}}^{\text{new}}$ and $T_{\text{{reply}}_{\text{1}}}^{\text{new}}$ are updated according to the selected $\Delta t$ and then continue to complete the ranging process. This approach results in a large difference between the arrival time of the legitimate Response message and the attack signal to the Initiator, thus avoiding the occurrence of the distance variation caused by the attack.
II-C Problem Statement
In the Ghost Peak attack, the timestamp estimated by the leading edge detection algorithm at the time offset $t$ is valid if the forged STS segment used by the attacker satisfies:
$\left|{\sum\limits_{i=0}^{\text{N}}{a_{\text{L}}[i]p(t)a_{\text{A}}[i-n]p^{*}(t)}}\right|>\max\left\{P_{\text{max}}\text{T}_{\text{m}},P_{\text{rms}}\text{T}_{\text{p}}\right\},$ (2)
where $t=~{}\frac{n}{\text{F}_{\text{{s}}}}\left(n=1,2,\cdots,\text{T}_{\text{{BTW}}}\text{F}_{\text{{s}}}\right)$ denotes the time corresponding to the sampling point, $\text{F}_{\text{{s}}}$ is the sampling rate, $\text{T}_{\text{{BTW}}}$ is the size of the BTW, $p(t)$is the pulse shape, $P_{\text{max}}$ and $P_{\text{rms}}$ are the maximum and root mean square power of the cross-correlation spectrum, $\text{T}_{\text{m}}$ and $\text{T}_{\text{p}}$ are the thresholds MPEP and PAPR corresponding to the leading edge detection algorithm, $a_{\text{L}}[n],a_{\text{A}}[n]\in\left\{-1,1\right\}$ are local STS template and the attacker’s STS, and N is the length of STS.
Since STS is a pseudo-random sequence modulated by Binary Phase Shift Keying (BPSK), it should satisfy:
${\mathbb{P}({a_{\text{L}}[i]a_{\text{A}}[i-n]=-1})}={\mathbb{P}({a_{\text{L}}[i]a_{\text{A}}[i-n]=1})}=0.5.$ (3)
Let $X\backsim\mathcal{B}(\text{N},0.5)$ be a random variable following binomial distribution, then the random variable $\sum\limits_{i=0}^{\text{N}}{a_{\text{L}}[i]~{}a_{\text{A}}[i-n]}$ follows:
${\sum\limits_{i=0}^{\text{N}}{a_{\text{L}}[i]~{}a_{\text{A}}[i-n]}}=2X-\text{N}.$ (4)
For simplicity, let $\max\left\{P_{\text{max}}\text{T}_{\text{m}},~{}P_{\text{rms}}\text{T}_{\text{p}}\right\}=\theta$, thus the attack success probability ${\mathbb{P}}_{\text{s}}(t)$ at the time offset $t$ is[15]:
${\mathbb{P}_{\text{s}}}(t)$ $=\left({\left|{2\mathcal{B}\left({\text{N},0.5}\right)-\text{N}}\right|\left|{p(t)}\right|^{2}>\theta}\right)$ (5)
$=2\left({2\mathcal{B}\left({\text{N},0.5}\right)-\text{N}~{}<\frac{-\theta}{\left|{p(t)}\right|^{2}}}\right)$
$=2F_{\text{BIN}}\left({0.5\text{N}-0.5\theta{(|{p(t)}|}^{2})^{-1}|{\text{N},0.5}}\right)$
$\leq 2exp\left({-\theta{(\left|{p(t)}\right|}^{2})^{-1}}\right),$
where $F_{\text{BIN}}(\cdot|\text{N},0.5)$ denotes the cumulative probability density function (CDF) of the binomial distribution $\mathcal{B}(\text{N},0.5)$, the upper bound is derived by Hoeffding’s inequality. It can be concluded that in addition to the thresholds, the power has a decisive influence on the attack success rate, which also explains why the power of the STS segment is extremely increased in the Ghost Peak attack.
Meanwhile, since UWB receivers usually use the cross-correlation result of the STS segment as the Channel Impulse Response (CIR) for data demodulation, in order to ensure that the Payload field can be demodulated correctly and does not affect the detection of the start of frame delimiter (SFD) field (as shown in Fig. 3), the time offset $t$ should satisfy:
${T_{\text{SFD}}<t<T_{\text{payload}}}$ (6)
where $T_{\text{SFD}}$ and $T_{\text{payload}}$ refer to the thresholds greater than the end instant of the SFD field and less than the beginning instant of the Payload field, respectively. These two thresholds ensure that the receiver is able to complete the reception of the ranging message, which is one of the preconditions for the success of the attack. We aim to reduce the attack success rate and improve the security of UWB ranging by breaking the requirement that the transmit time of the attack message and the legitimate message need to be synchronously aligned.
Figure 3: Packet Configuration 1 in IEEE 802.15.4z[5]
III Security Ranging Strategy Design
III-A Overall Procedure
Low complexity and low power consumption of the ranging algorithm are crucial considerations in the field of Internet of Things (IoT). To adapt to various scenarios, we propose a UWB secure ranging strategy utilizing random time-hopping mechanism, which mainly consists of an attack detection module and time-hopping DS-TWR as shown in Fig. 4.
Figure 4: A security ranging strategy with attack detection and time-hopping DS-TWR for UWB
The purpose of attack detection is to enable the UWB ranging system to recognize an ongoing attack, indicated by the output $S$ of the attack detection module being equal to 1. Then the ranging mode switches from the classical DS-TWR with lower power consumption to the time-hopping DS-TWR with higher power consumption but higher security. The principle of channel reciprocity based attack detection is to compress and quantize the CIR feature using an autoencoder in a complete ranging process, then transmit them using the Payload field, and finally detect the attack by comparing the CIR feature of both ranging sides. The specific procedure of attack detection is depicted in Fig. 4. It is not the focus of this paper and the detailed algorithm can be referred in[14].
As described in Sec. II-B, the security of the time-hopping ranging model is achieved by randomly varying the transmit time of the ranging messages. In brief, the random transmit time of legitimate messages makes it difficult for an attacker to accurately transmit the forged attack signal during the reception of the corresponding legitimate message, thus breaking the synchronization conditions required for the attacks.
The range of selectable random time-hopping delay $\Delta t$ is wide. In practical UWB applications, the Time of Flight (TOF) is generally in the order of nanosecond and the delay between each ranging message is in the order of millisecond. Basically, all kinds of distance attack methods must ensure that the attack signal and legitimate signal arrive at the receiver at a highly consistent time. Therefore, it is easy to select a set of $\Delta t$ values that can invalidate the attack without affecting the whole ranging process.
III-B Random Time-hopping Ranging without Redundant Signaling Overheads
In time-hopping DS-TWR, both ranging sides also need to obtain the specific random time-hopping delay value $\Delta t$ in real time, which will bring additional signaling overhead. In the generation process of the STS with AES encryption mechanism shown in Fig. 5, there exists a counter whose value is incremented each time a ranging message is transmitted or received and the counter values of both ranging sides remain consistent. Accordingly, this paper proposes an anti-attack time-hopping ranging scheme without redundant signaling consumption, which is implemented as follows:
III-B1 Random delay value preselection
Sort out the available $\Delta t$ values that invalidate the attack without affecting the entire ranging process and store them as a hash table to ensure quick access.
III-B2 STS counter value reading
When the Initiator and Responder transmit or receive a ranging message, read their respective counter values.
III-B3 Delay synchronization and distance calculation
Both ranging sides use their own counter value as a random seed to randomly select a $\Delta t$ value from the stored hash table. Then they complete the ranging process using the modified $T_{\text{round1}}^{\text{new}}$ and $T_{\text{reply1}}^{\text{new}}$.
Figure 5: STS generation procedure in IEEE 802.15.4z[5]
III-C Performance Analysis
Generally it can be assumed that $t$ follows the uniform distribution $\text{U}(\text{T}_{\text{min}},\text{T}_{\text{max}})$. Taking into consideration (5) and (6), the attack success probability ${\mathbb{P}_{\text{s}}}^{{}^{\prime}}(t)$ at the time offset $t$ becomes:
${\mathbb{P}_{\text{s}}}^{{}^{\prime}}(t)$ $={\mathbb{P}_{\text{s}}}(t)~{}\mathbb{P}({T_{\text{SFD}}<t<T_{\text{payload}}})$ (7)
$=2F_{\text{BIN}}({0.5\text{N}-0.5\theta{(|{p(t)}|}^{2})^{-1}|{\text{N},0.5}})\frac{\Delta t_{0}}{\Delta t_{1}},$
where $\Delta t_{0}=T_{\text{payload}}-T_{\text{SFD}}$ and $\Delta t_{1}=T_{\text{max}}-T_{\text{min}}$.
If the selected random hopping delay $\Delta t$ ranges between $T_{\text{min}}^{\text{hop}}$ and $T_{\text{max}}^{\text{hop}}$, $\Delta t$ can be considered to follow the uniform distribution $\text{U}(T_{\text{min}}^{\text{hop}},T_{\text{max}}^{\text{hop}})$ and be independent from $t$. Furthermore, to guarantee the success of the attack, $\Delta t$ and $t$ should satisfy:
$T_{\text{SFD}}<t-\Delta t<T_{\text{payload}}.$ (8)
Let the random variable $Y=t-\Delta t$, the attack success probability ${\mathbb{P}_{\text{s}}}^{{}^{\prime\prime}}(t)$ after adopting the random time-hopping mechanism becomes:
${\mathbb{P}_{\text{s}}}^{{}^{\prime\prime}}(t)$ $={\mathbb{P}_{\text{s}}}(t)~{}\mathbb{P}({T_{\text{SFD}}<t-\Delta t<T_{\text{payload}}})$ (9)
$=\left(2F_{\text{BIN}}({0.5\text{N}-0.5\theta{(|{p(t)}|}^{2})^{-1}|{\text{N},0.5}})\right.$
×$\left.\times{\int_{T_{\text{SFD}}}^{T_{\text{payload}}}f_{Y}}(y)dy\right),$
where ${f_{Y}}(y)$ is the probability density function (pdf) of the random variable $Y$, which is formed by the convolution of two uniform distributions.
Let $\Delta t_{2}=T_{\text{max}}^{\text{hop}}-T_{\text{min}}^{\text{hop}}$, then ${f_{Y}}(y)$ is an isosceles trapezoid between $T_{\text{max}}^{\text{hop}}$ and $T_{\text{min}}^{\text{hop}}$, with its height and upper base $\frac{1}{\Delta t_{2}}$ and $\Delta t_{2}-\Delta t_{1}$ respectively. Since the range of delay $\Delta t_{2}$ for the random time-hopping proposed in this paper is much larger than the original time interval $\Delta t_{1}$ (i.e. $\Delta t_{2}\gg\Delta t_{1}$), ${f_{Y}}(y)$ can be approximated as a rectangle with bottom $\Delta t_{2}$ and height $\frac{1}{\Delta t_{2}}$. Therefore, the reduction factor $G$ of the attack success probability (the gain of this strategy) is:
$G=\frac{\Delta t_{2}}{\Delta t_{1}}\rightarrow\infty$ (10)
On the other hand, the complexity of this security ranging strategy depends entirely on the attack detection module, while the random time-hopping ranging module does not increase the complexity at all.
IV Evaluation
IV-A Numerical Simulation Evaluation
The signal parameter configurations for simulation are detailed in Table I.
TABLE I: Configurations for legitimate and attack signals
Parameter Legitimate signals Attack signals
Mode BPRF BPRF
Preamble spreading factor 4 9
Preamble code index 9 9
SFD number 0 0
Modulation BPSK+BPM BPSK+BPM
Payload encoding RS &convolution RS &convolution
Samples of per pulse 4 4
Preamble duration 64 64
STS segment length 64 64
The processing flow of the receiver used in this paper includes shaping filtering, downsampling, cross-correlation, timestamp estimation and data demodulation utilizing a RAKE receiver architecture. The length of BTW is fixed at 400 samples (at a sampling rate of 2 GHz). MPEP and PAPR are set to 0.5 and 2, respectively. Moreover, the distance between two legitimate devices is 10 m and an attack is considered successful when the measurement result is less than 5 m. 20,000 DS-TWR simulations under the Ghost Peak attack are performed for each case.
SIR is used to denote the ratio between the power of the STS segment of the legitimate signal and that of the attack signal, and Signal-to-Noise Ratio (SNR) in the environment is set to -10 dB. The attack success probability simulated under different SIR and attack synchronization time $T_{\text{sy}}$ is shown in Table II.
TABLE II: Attack success probability (%) under different SIR and attack synchronization time $T_{\text{sy}}$
$T_{\text{sy}}$ SIR (dB)
(microsecond) -20 -22 -24 -26 -28 -30
-2.5 4% 4% 1.5% 1% 1% 1%
-2 4% 6% 4.5% 4% 1.5% 1%
-1.5 3.5% 4% 3.5% 2% 2.5% 1%
-1 20.5% 22.5% 29% 32.5% 21% 9%
-0.5 16% 22% 21% 24% 23% 22.5%
0 13% 14% 23% 25.5% 24.5% 19.5%
0.5 17% 23% 25% 24.5% 20.5% 19.5%
1 11.5% 20% 17% 18.5% 11.5% 6%
1.5 5% 0% 0.5% 0% 0% 0%
2 4.5% 0% 0% 0% 0% 0%
2.5 2% 0% 0% 0% 0% 0%
Figure 6: Missing detection probability and false alarm probability of the attack detection scheme
Table II shows that the attack success probability decreases significantly when the attack synchronization time exceeds a certain range because it affects the detection of the SFD field and the demodulation of the PHR field at the receiver. According to the IEEE 802.15.4z protocol, there is a 128-bit gap between the front and the end of the STS segment, which can attenuate the impact of the attack signal on the SFD field detection and PHR field demodulation. Moreover, the peak value generated by the attack signal in the cross-correlation spectrum will be lower than the thresholds of the leading edge detection algorithm if the power of the STS segment of the attack signal is too low and the strongest path obtained by the leading edge detection algorithm may be changed if the power of the STS segment of the attack signal is too high, and attacks will fail in either of these circumstances.
To evaluate the performance of the proposed attack detection scheme in DS-TWR, we simulate under the Ghost Peak attack condition to obtain the missing detection probability and false alarm probability of the attack detection scheme with respect to SNR. As shown in Fig. 6, the attack detection scheme has a relatively low probability of missing detection and false alarm, and is robust to SNR. The reliability of the attack detection scheme provides a solid foundation for applying the random time-hopping DS-TWR.
The variation curve of attack success probability with SNR before and after using the classical DS-TWR and time-hopping DS-TWR under different SNR and SIR is shown in Fig. 7, where the random time-hopping delay $\Delta t$ is set to be from 15 to 20 microsecond.
Figure 7: Attack success probability before and after adopting time-hopping DS-TWR
Fig. 7 indicates that the time-hopping DS-TWR can reduce the attack success probability to 0% in the simulation, reaching the order of one in ten thousand regardless of the attack success probability under normal circumstances. The proposed security strategy is robust to both SIR and SNR, verifying that the anti-attack mechanism performs well in resisting attacks.
IV-B Experimental Validation
Figure 8: Practical test environment
We validate the proposed security ranging strategy based on random time-hopping mechanism using commercial ranging devices embedded with Qorvo DW3110 chip in the Ghost Peak attack scenario (an indoor 10 m environment shown in Fig. 8). The results of the attack success probability are shown in Table. III with $\Delta t$ set in the range of -2 to 3600 microsecond in the Mode1 mode and the more vulnerable SDC mode of STS[16].
TABLE III: Attack success probability of practical validation
STS mode Befor anti-attack After anti-attack
SDC 60.067% 0%
Mode1 0.3% 0%
According to listed results, the measurement of the chips can remain stable for tens of hours after adopting the proposed secure ranging strategy without redundant signaling overhead, which indicates that the proposed strategy performs satisfactorily in the practical environment.
V Conclusion
In order to improve the security performance of existing UWB ranging, this paper proposes the time-hopping DS-TWR which is able to significantly reduce the attack success probability of various types of distance attacks represented by the Ghost Peak attack by increasing the randomness of the time of transmitting legitimate messages in the ranging process. Furthermore, combined with the attack detection scheme, the proposed secure ranging strategy is able to adapt to various UWB ranging application scenarios with very low complexity. However, this scheme mainly increases the randomness of the transmit time of the ranging message, and future work will incorporate more aspects (e.g. power and frame format) to optimize the security performance of the proposed strategy.